OpenStudy (anonymous):

Suppose $12,000 is invested into an account where interest is compounded twice a year. After 7 years, the balance is $13,793.70. What was the interest rate as a percent?

OpenStudy (anonymous):

\[P\left(1+\frac{R}{N}\right)^{\text{NT}}=A\]

OpenStudy (anonymous):

We know P=12000 N= 2 (compounded twice a year) T=7 years A=13793.70 (final balance) R=?

OpenStudy (anonymous):

hold on let me try it and u check alright

OpenStudy (anonymous):

o.o47

OpenStudy (anonymous):

Did that check out?

OpenStudy (anonymous):

that is it i think wht u mean check out?

OpenStudy (anonymous):

I got different answer\[12000\left(1+\frac{R}{2}\right)^{14}\text{=}13793.70\] \[ \left(1+\frac{R}{2}\right)^{14}\text{==}\frac{13793.70}{12000}\] \[\left(1+\frac{R}{2}\right)^{14}==1.14948\] \[\left(\left(1+\frac{R}{2}\right)^{14}\right)^{\frac{1}{14}}==(1.14948)^{\frac{1}{14}}\] \[1+\frac{R}{2}==1.01\] \[\frac{R}{2}==.01\] R=.02